40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Apr 6 02:07:40 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
    Reply  Wed Apr 10 00:33:09 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source Modulation_frequency_100Hz.jpgModulation_frequency_30Hz.jpgModified_setup.JPG
       Reply  Wed Apr 10 16:58:54 2019, rana, Update, IOO, fiber MZ for NPRO freq noise measurements 
          Reply  Wed Apr 10 22:59:22 2019, gautam, Update, IOO, Spooled fiber 
             Reply  Thu Apr 11 09:05:06 2019, Anjali, Update, IOO, Spooled fiber Frequecy_modulation_32_Hz.pdfFrequecy_modulation_140_Hz.pdfFrequecy_modulation_300_Hz.pdfWithout_modulation.pdfNew_fiber_spool.JPG
       Reply  Fri Apr 12 01:22:27 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
          Reply  Mon Apr 15 22:39:10 2019, gautam, Update, Frequency noise measurement, Alternate setup with PSL pickoff 
             Reply  Wed Apr 17 00:43:38 2019, gautam, Update, Frequency noise measurement, MZ interferometer ---> DAQ 
                Reply  Thu Apr 25 03:32:25 2019, Anjali, Update, Frequency noise measurement, MZ interferometer ---> DAQ Time_domain_output.pdfFrequency_noise.pdfschematic_heterodyne_setup.pngNoise_budget_1_micron_in_Hz_per_rtHz.pdf
                   Reply  Thu Apr 25 10:25:19 2019, gautam, Update, Frequency noise measurement, Homodyne v Heterodyne 
                      Reply  Thu Apr 25 15:47:54 2019, Anjali, Update, Frequency noise measurement, Homodyne v Heterodyne Experimental_setup.JPG
Message ID: 14520     Entry time: Sat Apr 6 02:07:40 2019     Reply to this: 14529
Author: Anjali 
Type: Update 
Category: Frequency noise measurement 
Subject: Frequency noise measurement of 1 micron source 
  • The alignment of the output beam from the delayed path of MZI to the photodetector was disturbed when we did the polarisation characterisation yesterday. So, today we tried to align the output beam from the delayed path of MZI to the detector .
  • We then observed the beat output from the detector on oscilloscope.We initialy observed a dc shift . We then applied a frequency modulation on the input laser and observed the output on oscilloscope. We expected to see variation in output frequency in accordance with variation of input frequency modulation. But we didnt observe this and we were not really getting the interference pattern. 
  • We tried to make the alignment better. With a better alignment, we could see the interference pattern. We also observed that the output frequency was varying in accordance with variation in the input frequency modulation. We would expect a better result with proper mode matching of the two beams on the photodetector.
ELOG V3.1.3-